# Isolated one-phonon mixed-symmetry state of the radioactive neutron-rich nuclide 132Te

### v1

@@ Acknowledgements.
The experiment was carried out in 2021 during the peak of the COVID-19 pandemic. For that reason, travel to the experiment was not possible and sincere thanks are due to the colleagues at IFIN-HH who performed all hands-on work during the experiment. Also the staff of the IFIN-HH tandem accelerator is acknowledged for providing optimal beam conditions. This work was supported by the BMBF under Grant No. 05P21RDCI2, by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – Project-ID 499256822 – GRK 2891 ’Nuclear Photonics’, and by Project ELI-RO/DFG/2023_001 ARNPhot funded by the Institute of Atomic Physics, Romania. G.R. acknowledges the support by the Bulgarian Ministry of Education and Science, within the National Roadmap for Research Infrastructures (object CERN) and T.O. support by the German Alexander von Humboldt-Foundation.References

### v2

@@ Acknowledgements.
The experiment was carried out in 2021 during the peak of the COVID-19 pandemic. For that reason, travel to the experiment was not possible and sincere thanks are due to the colleagues at IFIN-HH who professionally performed all hands-on work during the experiment. Also the staff of the IFIN-HH tandem accelerator is acknowledged for providing optimal beam conditions and for target preparation. This work was supported by the German BMBF under Grant No. 05P21RDCI2, by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – Project-ID 499256822 – GRK 2891 ’Nuclear Photonics’, by Project ELI-RO/DFG/2023_001 ARNPhot funded by the Institute of Atomic Physics, Romania, and by the Romanian Nucleu project No. PN 23 21 01 02. K.G., D.K. and G.R. acknowledge support by the European Union-NextGenerationEU, through the National Recovery and Resilience Plan of the Republic of Bulgaria, project No. BG-RRP-2.004-0008-C01, and by the Bulgarian Ministry of Education and Science, within the National Roadmap for Research Infrastructures (object CERN) and T.O. support by the German Alexander von Humboldt-Foundation. The data that support the findings of this article are not publicly available. The data are available from the authors upon reasonable request.References
@@
- Heyde and Sau [1986] K. Heyde and J. Sau, Phys. Rev. C 33, 1050 (1986).
- Covelle et al. [2007] A. Covelle, L. Coraggio, A. Gargano, and N. Itaco, Prog. Part. Nucl. Phys. 59, 401 (2007).
- Stuchbery and Wood [2022] A. E. Stuchbery and J. L. Wood, To shell model, or not to shell model, that is the question, Physics 4, 697 (2022).
@@
- Moschner et al. [2016] K. Moschner, A. Blazhev, J. Jolie, N. Warr, P. Boutachkov, P. Bednarczyk, K. Sieja, A. Algora, F. Ameil, M. A. Bentley, S. Brambilla, N. Braun, F. Camera, J. Cederkäll, A. Corsi, M. Danchev, D. DiJulio, C. Fahlander, J. Gerl, A. Giaz, P. Golubev, M. Górska, J. Grebosz, T. Habermann, M. Hackstein, R. Hoischen, I. Kojouharov, N. Kurz, N. Mărginean, E. Merchán, T. Möller, F. Naqvi, B. S. Nara Singh, C. Nociforo, N. Pietralla, S. Pietri, Zs.. Podolyák, A. Prochazka, M. Reese, P. Reiter, M. Rudigier, D. Rudolph, T. Sava, H. Schaffner, L. Scruton, J. Taprogge, T. Thomas, H. Weick, A. Wendt, O. Wieland, and H.-J. Wollersheim, Phys. Rev. C 94, 054323 (2016).
- Werner et al. [2008] V. Werner, N. Benczer-Koller, G. Kumbartzki, J. D. Holt, P. Boutachkov, E. Stefanova, M. Perry, N. Pietralla, H. Ai, K. Aleksandrova, G. Anderson, R. B. Cakirli, R. J. Casperson, R. F. Casten, M. Chamberlain, C. Copos, B. Darakchieva, S. Eckel, M. Evtimova, C. R. Fitzpatrick, A. B. Garnsworthy, G. Gürdal, A. Heinz, D. Kovacheva, C. Lambie-Hanson, X. Liang, P. Manchev, E. A. McCutchan, D. A. Meyer, J. Qian, A. Schmidt, N. J. Thompson, E. Williams, and R. Winkler, Phys. Rev. C 78, 031301 (2008).
- Casperson et al. [2013] R. Casperson, V. Werner, and S. Heinze, Physics Letters B 721, 51 (2013).
@@
- Brown et al. [2005] B. A. Brown, N. J. Stone, J. R. Stone, I. S. Towner, and M. Hjorth-Jensen, Phys. Rev. C 71, 044317 (2005).
@@
- Shimizu [2013] N. Shimizu, (2013), arXiv:1310.5431 [nucl-th] .
- Schimizu et al. [2019] N. Schimizu, T. Mizusaki, Y. Utsuno, and Y. Tsunoda, Comp. Phys. Comm. 244, 372 (2019).
- Gray et al. [2020] T. J. Gray, J. M. Allmond, A. E. Stuchbery, C.-H. Yu, C. Baktash, A. Gargano, A. Galindo-Uribarri, D. C. Radford, J. C. Batchelder, J. R. Beene, C. R. Bingham, L. Coraggio, A. Covello, M. Danchev, C. J. Gross, P. A. Hausladen, N. Itaco, W. Krolas, J. F. Liang, E. Padilla-Rodal, J. Pavan, D. W. Stracener, and R. L. Varner, Phys. Rev. Lett. 124, 032502 (2020).
- Hicks et al. [2022] S. F. Hicks, A. E. Stuchbery, T. H. Churchill, D. Bandyopadhyay, C. B. R., B. J. Coombes, and C. M. Davoren, Phys. Rev. C 105, 024329 (2022).